1. During a basketball game, a player complains that the ball isn’t inflated enough. To meet international basketball standards a ball dropped from 1.8 m should bounce to a height of 1.2 to 1.4 m. The ball only rebounded 1.1 m indicating that the ball needed more air. What was the coefficient of restitution of the basketball when it rebounded to a height of 1.1 m?
Answer #1

velocity of object before bouncing

u^2 = 2 gh

u^2 = 2* 9.8* 1.8

u = 5.94 m/s

Velocity of ball after bouncing

v^2 = 2* 9.8* 1.1

v = 4.643 m/s

now, COR is given by

COR = v/ u = 4.643 / 5.94

COR = 0.7816

========

do comment in case any doubt, will reply for sure. Goodluck